# Deep Learning for Subspace Regression

### v2

@@ 4.1 Eigenspace prediction
We considered several eigenvalue problems: (i) $D=1$ eigenproblem with Schrödinger operator (3) with parametric family of expanded Morse oscillator [20], (ii) $D=2$ Schrödinger operator (3) with parametric family of two expanded Morse oscillators [9], (iii) $D=2$ two datasets, $k_{1}=k_{2}$ and $k_{1}\neq k_{2}$, for elliptic eigenproblem (5) (left-hand side of the equation) with contrast coefficient sampled from gaussian random field, (iv) $D=3$ dataset for elliptic eigenproblem with diffusion coefficient $k_{1}=k_{2}$ As the main neural network architecture we used FFNO [31], a descendant of FNO [21], and performed extensive grid searches for all experiments. Details on dataset generation and training protocol are available in Appendix F. To contextualise subspace regression we provide results for two baselines.
@@
Loss without QRQR scales much better for larger subspace sizes. In Figure 1(a) we demonstrate wall clock training time for $L_{1}(A,B)$ and $L_{2}(A,B;z)$ (least squares problem is solved with normal equation) per epoch on the same hardware for the same FFNO architectures. For small subspace sizes the training time the methods are roughly on par, but with the increase of subspace size QR starts to drastically slow down training with $L_{1}$ loss.
@@
Loss $L_{2}(A,B;z)$ may become unstable. From results summarised in Figure 1(a) one can assume that $L_{2}(A,B;z)$ is always preferable. Results for $D=3$ problem (elliptic eigenproblem, grid $30\times 30\times 30$, prediction of first 33 eigenvectors) summarised in Table 2 reveal a more nuanced picture. Loss $L_{2}(A,B;z)$ clearly performs worse than $L_{1}(A,B)$ and even fails for subspace size $24$. The reason for that is numerical instability of solvers based on the normal equation. To stabilize $L_{2}(A,B;z)$ we apply Cholesky-QR2 [37]. The results for stabilised loss shows that accuracy becomes comparable to $L_{1}(A,B)$ and even slightly better for larger subspace dimensions.
@@ 4.2 Parametric PDE problems
We considered two PDEs: (i) $D=1+1$ viscous Burgers equation (related to benchmark from [21]), (ii) $D=2$ elliptic problems (5). Our main operator is FFNO and the solutions strategy we use is classical intrusive POD [^2] Recall, that when coefficients in the reduced basis expansion are predicted by some model we have non-intrusive POD. When basis is used to generate reduced ODE that is later integrated we have intrusive POD.. For datasets description and training details see Appendix G. We compare subspace regression with several methods.
@@
4.3 Iterative methods for linear systems
Seemingly minor variations in problem setting can lead to substantial variations in the complexity of the learning problem. Initially for the Jacobi method we posed a subspace regression problem as an approximation of leading eigenspaces of error propagation matrix $I-D^{-1}A$, where $D$ is diagonal of AA. Neural networks with and without subspace embedding completely failed to learn. After inspection of the dataset we found that the leading eigenspace contains a complicated mixture of functions with low and high frequencies. Since the learning problem appeared to be completely hopeless, we reformulated subspace regression using error propagation matrix of damped Jacobi iteration $I-\omega D^{-1}A$ with $\omega=0.9$. In contrast to the standard Jacobi method, the damped version leads to subspaces formed by low frequency functions. As evident from Figure 3(b) the resulting mapping is easily learnable. A more detailed report can be found in Appendix H.
@@ Appendix A Proof of Theorem 1
    where $P_{A}=A\left(A^{\top}A\right)^{-1}A^{\top},P_{B}=B\left(B^{\top}B\right)^{-1}B^{\top}$ are orthogonal projectors on the columns spaces of AA and $B$. When QR decompositions $A=Q_{A}R_{A},B=Q_{B}R_{B}$ are available, projectors become $P_{A}=Q_{A}Q_{A}^{\top},P_{B}=Q_{B}Q_{B}^{\top}$ and identity (8) can be verified by algebraic manipulations
@@
    where $P_{A}=A\left(A^{\top}A\right)^{-1}A^{\top}$ is orthogonal projector on the columns space of AA. Using $I=\left(I-P_{A}\right)+P_{A}$, and $A\left(I-P_{A}\right)=\left(I-P_{A}\right)A=0$ we obtain
@@
    The last equality holds since $P_{A}Q_{B}$ and AA share the same columns space. Given that $P_{A}$ does not depend on representative AA from $\lceil A\rceil$, and that $L_{2}(A,B;z)$ depends on AA only via $P_{A}$, we conclude that the same is true for $L_{2}(A,B;z)$.
@@ Appendix B Proof of Theorem 2
In most parts of the text we assumed working with the non-compact Stiefel manifold and in this theorem we have data on the compact Stiefel manifold (see [1] for definitions). We specify how one can compute $Q_{A}$ and $\dot{Q}_{A}$ having AA and $\dot{A}$. One may start from any stable version of Cholesky QR, e.g., [14], [37], and obtain
@@
where $y_{i}$ are columns of $Y$, $u_{i}$ are columns of $U$, $\sigma_{i}$ are diagonal elements of $\Sigma$, $\widetilde{\Sigma}=\Sigma-(\sigma_{1}-1)e_{1}e_{1}^{\top}$, that is, $\widetilde{\Sigma}$ can be obtained from $\Sigma$ by replacing $\Sigma_{11}=\sigma_{1}$ by 11. Clearly $B(t)$ is geodesic and $\left\|\dot{B}(t)\right\|_{F}^{2}=\left\|\dot{A}(t)\right\|_{F}^{2}-\sigma_{1}<\left\|\dot{A}(t)\right\|_{F}^{2}$. Next we show that principal angles between $A(t)$ and $B(t)$ are all zero. To see this we observe that
@@ Appendix D Proof of Theorem 3
To illustrate this process, consider $E(z_{1},\dots,z_{D})=a_{1}z_{1}^{2}+a_{2}z_{2}^{2}$, where $a_{2}\gg a_{1}$. If we follow procedure outlined above we will see that first lattice points encountered are $(1,1),(2,1),(3,1),(4,1),\dots$. So for considered $a_{1},a_{2}$ the subspace of first 33 eigenvectors is a span of $\phi_{1,1},\phi_{2,1},\phi_{3,1}$, and the eigenvector that appears on position 33 is $\phi_{3}$. To describe the map from $a_{1},\dots,a_{D}$ to $\phi_{k}$ or $V_{k}$, this procedure needs to be repeated for all possible positive values of real coefficients $a_{1},\dots,a_{D}$.
From the algorithm above one can deduce that for given $a_{1},\dots,a_{D}$ the first time eigenvector with indices $i_{1},\dots,i_{D}$ appears in the sequence of eigenvectors is the first time ellipsoid crosses $i_{1},\dots,i_{D}$. The position $k$ of this eigenvector will be proportional to the normalised volume of the ellipsoid $V_{e}(a_{1},\dots,a_{D})\big/V_{s}$, where $V_{e}(a_{1},\dots,a_{D})$ is a volume of $D$ dimensional ellipsoid with semi-axes $a_{1},\dots,a_{D}$ and $V_{s}$ is a volume of $D$ dimensional sphere with radius 11.
@@ Appendix E Monte Carlo experiments for eigenproblem with constant coefficients
To illustrate consequences of Theorem 3 we perform simple Monte Carlo experiment. For the case $D=3$, we generate $a_{1},a_{2},a_{3}$ from uniform distribution on $[0,1]$ repeatedly and record distinct subspaces of dimension $k=10$. We select unique colour for each subspace and draw them for each point $a_{1},a_{2},a_{3}$ on the plane perpendicular to $\begin{pmatrix}1&1&1\end{pmatrix}^{\top}$. This illustration appears in Figure 6(a). Theorem 3 suggest that the large number of distinct subspaces is a result of selection of $k$ vectors among a small number of candidates. This suggests we may decrease complexity of the function from coefficients to subspaces by predicting excessive number of vectors. In our experiments with neural networks this redundant mapping is learned, here we build the mapping with simple greedy strategy. In place of function $a_{1},a_{2},a_{3}\rightarrow V_{10}$ we consider $a_{1},a_{2},a_{3}\rightarrow V_{10}\cup\left\{v_{1},\dots,v_{k}\right\}$ where $v_{1},\dots,v_{k}$ are first $k$ most abundant eigenvectors. By appending additional vectors we decrease the number of distinct subspaces. The result of this greedy simplification appear in Figure 6(b) with $10$ additional vectors and in Figure 6(c) with 2020 additional vectors. As we see the number of distinct subspaces rapidly decreasing.
@@ F.1 Datasets
where $u,v$ are component of random normalised vector, $c=\sqrt{2}r_{e}$. Potentials $V_{1}$ and $V_{2}$ are i.i.d. with parameters: order of polynomial is 22, $r_{e}$ is uniformly distributed on $[1,5]$, $d$ is uniformly distributed on $[10,40]$, coefficients of $\widetilde{q}$ are sampled from uniform distribution on $[0,3]$ and $c$ is sampled from uniform distribution on $[10,13]$. To discretise the problem we use finite difference and uniform $100\times 100$ grid on $7,7]^{2}$.
@@ F.2 Architectures and training
For $D=3$ grid search is not practical, so we select $N_{\text{layers}}=4$, $N_{\text{features}}=128$, $N_{\text{modes}}=16$, $\text{lr}=10^{-3}$, $\gamma_{\text{decay}}=0.5$, $N_{\text{decay}}=100$. The size of the train set is $800$, the size of test set is $200$. The number of epochs is $1000$. Architecture is training to approximate subspace spanned by first 33 eigenvectors.
F.3 Additional results

### v3

@@ 4.1 Eigenspace prediction
We considered several eigenvalue problems: (i) $D=1$ eigenproblem with Schrödinger operator (3) with parametric family of expanded Morse oscillator [21], (ii) $D=2$ Schrödinger operator (3) with parametric family of two expanded Morse oscillators [9], (iii) $D=2$ two datasets, $k_{1}=k_{2}$ and $k_{1}\neq k_{2}$, for elliptic eigenproblem (5) (left-hand side of the equation) with contrast coefficient sampled from gaussian random field, (iv) $D=3$ dataset for elliptic eigenproblem with diffusion coefficient $k_{1}=k_{2}$. In all experiments we used FFNO architecture [32] , a modification of FNO [22], and performed extensive hyperparameter grid search. Details on dataset generation and training protocol are available in Appendix F. To contextualise subspace regression we provide results for two baselines.
@@
Loss without $QR$ scales much better for larger subspace sizes. In Figure 1(a) we demonstrate wall clock training time for $L_{1}(A,B)$ and $L_{2}(A,B;z)$ (least squares problem is solved with normal equation) per epoch on the same hardware for the same FFNO architectures. For small subspace sizes the training time the methods are roughly on par, but with the increase of subspace size QR starts to drastically slow down training with $L_{1}$ loss.
@@
Neural networks trained with subspace embedding technique learn smoother maps. Theorem 2 suggests it is possible to decrease derivative by embedding of geodesics into a larger space. This provides only a circumstantial evidence that the same may happen when neural networks are trained with subspace embedding technique. In Appendix F we gather empirical results that support such conclusion. The results are based on several “smoothness indicators”: the error of linear model, Frobenius norm of derivative, and mean cosine of angles between subspaces at nearby points. We refer interested readers to Appendix F.5.
Subspace regression can speed-up classical iterative eigensolvers. As an example of hybrid approach we consider combination of subspace regression and LOBPCG [19]. LOBPCG is a classical iterative matrix-free eigensolver that can approximate extremal eigenspaces. To apply it in combination wit hsubspace regression we use trained neural network to predict subspace, and initialise LOBPCG with solution of reduced eigenproblem. Note, that the cost of such initialisation is negligible small comparing to the full cost of LOBPCG iterations. We observe $2$ to $3$ times faster converges and $2$ orders lower relative error on average when subspace regression is used for initialisation. More details are available in Appendix F.6.
@@
Loss $L_{2}(A,B;z)$ may become unstable. From results summarised in Figure 1(a) one can assume that $L_{2}(A,B;z)$ is always preferable. Results for $D=3$ problem (elliptic eigenproblem, grid $30\times 30\times 30$, prediction of first $3$ eigenvectors) summarised in Table 2 reveal a more nuanced picture. Loss $L_{2}(A,B;z)$ clearly performs worse than $L_{1}(A,B)$ and even fails for subspace size $24$. The reason for that is numerical instability of solvers based on the normal equation. To stabilize $L_{2}(A,B;z)$ we apply Cholesky-QR2 [38]. The results for stabilised loss shows that accuracy becomes comparable to $L_{1}(A,B)$ and even slightly better for larger subspace dimensions.
@@ 4.2 Parametric PDE problems
We considered two PDEs: (i) $D=1+1$ viscous Burgers equation, related to benchmark from [22]; (ii) $D=2$ elliptic problems (5). Our main operator is FFNO and the solutions strategy we use is classical intrusive POD [^4] Recall, that when coefficients in the reduced basis expansion are predicted by some model we have non-intrusive POD. When basis is used to generate reduced ODE that is later integrated we have intrusive POD.. For datasets description and training details see Appendix G. We compare subspace regression with several methods.
@@
4.3 Iterative methods for linear systems
Seemingly minor variations in problem setting can lead to substantial variations in the complexity of the learning problem. Initially for the Jacobi method we posed a subspace regression problem as an approximation of leading eigenspaces of error propagation matrix $I-D^{-1}A$, where $D$ is diagonal of $A$. Neural networks with and without subspace embedding completely failed to learn. After inspection of the dataset we found that the leading eigenspace contains a complicated mixture of functions with low and high frequencies. Since the learning problem appeared to be completely hopeless, we reformulated subspace regression using error propagation matrix of damped Jacobi iteration $I-\omega D^{-1}A$ with $\omega=0.9$. In contrast to the standard Jacobi method, the damped version leads to subspaces formed by low frequency functions. As evident from Figure 3(b) the resulting mapping is easily learnable. A more detailed report can be found in Appendix H.
@@ 5 Conclusion
6 Reproducibility statement
Code used for training, evaluation, and dataset generation is available on https://github.com/VLSF/subreg. In the current version detailed description of architectures, hyperparameters, dataset generation and training details appear in Appendix I, Appendix H, Appendix G, Appendix F.
@@ Appendix A Proof of Theorem 1
    where $P_{A}=A\left(A^{\top}A\right)^{-1}A^{\top},P_{B}=B\left(B^{\top}B\right)^{-1}B^{\top}$ are orthogonal projectors on the columns spaces of $A$ and $B$. When QR decompositions $A=Q_{A}R_{A},B=Q_{B}R_{B}$ are available, projectors become $P_{A}=Q_{A}Q_{A}^{\top},P_{B}=Q_{B}Q_{B}^{\top}$ and identity (8) can be verified by algebraic manipulations
@@
    where $P_{A}=A\left(A^{\top}A\right)^{-1}A^{\top}$ is orthogonal projector on the columns space of $A$. Using $I=\left(I-P_{A}\right)+P_{A}$, and $A\left(I-P_{A}\right)=\left(I-P_{A}\right)A=0$ we obtain
@@
    The last equality holds since $P_{A}Q_{B}$ and $A$ share the same columns space. Given that $P_{A}$ does not depend on representative $A$ from $\lceil A\rceil$, and that $L_{2}(A,B;z)$ depends on $A$ only via $P_{A}$, we conclude that the same is true for $L_{2}(A,B;z)$.
@@ Appendix B Proof of Theorem 2
In most parts of the text we assumed working with the non-compact Stiefel manifold and in this theorem we have data on the compact Stiefel manifold (see [1] for definitions). We specify how one can compute $Q_{A}$ and $\dot{Q}_{A}$ having $A$ and $\dot{A}$. One may start from any stable version of Cholesky QR, e.g., [14], [38], and obtain
@@
where $y_{i}$ are columns of $Y$, $u_{i}$ are columns of $U$, $\sigma_{i}$ are diagonal elements of $\Sigma$, $\widetilde{\Sigma}=\Sigma-(\sigma_{1}-1)e_{1}e_{1}^{\top}$, that is, $\widetilde{\Sigma}$ can be obtained from $\Sigma$ by replacing $\Sigma_{11}=\sigma_{1}$ by $1$. Clearly $B(t)$ is geodesic and $\left\|\dot{B}(t)\right\|_{F}^{2}=\left\|\dot{A}(t)\right\|_{F}^{2}-\sigma_{1}<\left\|\dot{A}(t)\right\|_{F}^{2}$. Next we show that principal angles between $A(t)$ and $B(t)$ are all zero. To see this we observe that
@@ Appendix D Proof of Theorem 3
In the proof we will write $F$ ad $G$ in place of $F_{k}$ and $G_{k}$ assuming that $k$ is fixed and the value of $k$ is evident from the context.
@@
To illustrate this process, consider $E(z_{1},\dots,z_{D})=a_{1}z_{1}^{2}+a_{2}z_{2}^{2}$, where $a_{2}\gg a_{1}$. If we follow procedure outlined above we will see that first lattice points encountered are $(1,1),(2,1),(3,1),(4,1),\dots$. So for considered $a_{1},a_{2}$ the subspace of first $3$ eigenvectors is a span of $\phi_{1,1},\phi_{2,1},\phi_{3,1}$, and the eigenvector that appears on position $3$ is $\phi_{3}$. To describe the map from $a_{1},\dots,a_{D}$ to $\phi_{k}$ or $V_{k}$, this procedure needs to be repeated for all possible positive values of real coefficients $a_{1},\dots,a_{D}$.
From the algorithm above one can deduce that for given $a_{1},\dots,a_{D}$ the first time eigenvector with indices $i_{1},\dots,i_{D}$ appears in the sequence of eigenvectors is the first time ellipsoid crosses $i_{1},\dots,i_{D}$. The position $k$ of this eigenvector will be proportional to the normalised volume of the ellipsoid $V_{e}(a_{1},\dots,a_{D})\big/V_{s}$, where $V_{e}(a_{1},\dots,a_{D})$ is a volume of $D$ dimensional ellipsoid with semi-axes $a_{1},\dots,a_{D}$ and $V_{s}$ is a volume of $D$ dimensional sphere with radius $1$.
@@ Appendix E Monte Carlo experiments for eigenproblem with constant coefficients
To illustrate consequences of Theorem 3 we perform simple Monte Carlo experiment. For the case $D=3$, we generate $a_{1},a_{2},a_{3}$ from uniform distribution on $[0,1]$ repeatedly and record distinct subspaces of dimension $k=10$. We select unique colour for each subspace and draw them for each point $a_{1},a_{2},a_{3}$ on the plane perpendicular to $\begin{pmatrix}1&1&1\end{pmatrix}^{\top}$. This illustration appears in Figure 6(a). Theorem 3 suggest that the large number of distinct subspaces is a result of selection of $k$ vectors among a small number of candidates. This suggests we may decrease complexity of the function from coefficients to subspaces by predicting excessive number of vectors. In our experiments with neural networks this redundant mapping is learned, here we build the mapping with simple greedy strategy. In place of function $a_{1},a_{2},a_{3}\rightarrow V_{10}$ we consider $a_{1},a_{2},a_{3}\rightarrow V_{10}\cup\left\{v_{1},\dots,v_{k}\right\}$ where $v_{1},\dots,v_{k}$ are first $k$ most abundant eigenvectors. By appending additional vectors we decrease the number of distinct subspaces. The result of this greedy simplification appear in Figure 6(b) with $10$ additional vectors and in Figure 6(c) with $20$ additional vectors. As we see the number of distinct subspaces rapidly decreasing.
@@ F.1 Datasets
where $u,v$ are component of random normalised vector, $c=\sqrt{2}r_{e}$. Potentials $V_{1}$ and $V_{2}$ are i.i.d. with parameters: order of polynomial is $2$, $r_{e}$ is uniformly distributed on $[1,5]$, $d$ is uniformly distributed on $[10,40]$, coefficients of $\widetilde{q}$ are sampled from uniform distribution on $[0,3]$ and $c$ is sampled from uniform distribution on $[10,13]$. To discretise the problem we use finite difference and uniform $100\times 100$ grid on $7,7]^{2}$.
@@ F.2 Architectures and training
Since architectures used are neural operators, we were able to speed up grid search by a factor of $7$ by first training architecture on coarse grid $32\times 32$, and then retraining $3$ most accurate architectures (according to the results of grid search on the coarse grid) of the fine $100\times 100$ grid.
For $D=3$ grid search is not practical, so we select $N_{\text{layers}}=4$, $N_{\text{features}}=128$, $N_{\text{modes}}=16$, $\text{lr}=10^{-3}$, $\gamma_{\text{decay}}=0.5$, $N_{\text{decay}}=100$. The size of the train set is $800$, the size of test set is $200$. The number of epochs is $1000$. Architecture is training to approximate subspace spanned by first $3$ eigenvectors.
@@
F.3 Additional results for elliptic eigenproblems
F.4 Generalisation to different grid size
For all subspace regression problems we use FFNO. Since FFNO is neural operator it should be discretisation agnostic. Here we report results for model trained on grid $100$ on $D=1$ quantum mechanics eigenproblem and tested on grids of higher resolution. For each grid we generated new test set from the same distribution as specified in Appendix F.1.
The results are available in Table 6. We see approximately linear increase of relative error with resolution. Owning to good initial accuracy on grid $N_{x}=100$ the relative error remains under $10\%$ for grid with $N_{x}=500$.
Table 6: Network is trained on resolution $N_{x}=100$ for $D=1$ QM problem and evaluated on grids with increased resolution.
$N_{x}$ $100$ $150$ $200$ $250$ $300$ $350$ $400$ $450$ $500$
test error, $\%$ $0.83$ $1.37$ $2.12$ $2.95$ $3.96$ $4.32$ $5.33$ $5.32$ $6.79$
F.5 Smoothness of neural networks trained with subspace embedding technique
To empirically measure smoothness of learned map, we introduce several “smoothness indicators”:
  1. 1.
    Taylor indicator
    $T[f_{1},f_{2};l]=\frac{\left\|\mathcal{N}(f_{1}+lf_{2})-\mathcal{N}(f_{1})-\left.\frac{d}{dl}\mathcal{N}(f_{1}+lf_{2})\right|_{l=0}l\right\|_{2}}{\left\|\mathcal{N}(f_{1}+lf_{2})\right\|_{2}}.$ (49)
    Taylor indicator is a relative error of linear model. One expect that: (i) unless $\mathcal{N}$ is a linear function, when $l$ increases relative error also increases; (ii) when smoothness increases Taylor indicator decreases.
  2. 2.
    Average cosine
    $C[f_{1},f_{2};l]=\frac{1}{D}\sum_{i=1}^{D}\cos_{i}(\mathcal{N}(f_{1}+lf_{2}),t_{1}).$ (50)
    where $\cos_{i}(A,B)$ are cosines of principle angles [8] and $t_{1}$ is target at point $f_{1}$. For smoother maps average cosine increases until it reaches maximal value of $1$.
  3. 3.
    Frobenius norm of the directional derivative
    $F[f_{1},f_{2};l]=\frac{1}{D}\left\|\left.\frac{d}{dl}\mathcal{N}(f_{1}+lf_{2})\right|_{l=0}\right\|_{F},$ (51)
    where $D$ is the subspace size. The magnitude of the directional derivative is computed by automatic differentiation and it is expected to decrease when smoothness increases.
Table 7: Frobenius norm of the directional derivative smoothness indicator (51).
$N_{\text{subspace}}$ $k_{1}=k_{2}$ $k_{1}\neq k_{2}$
$10$ $10.73$ $8.26$
2020 $9.72$ $8.15$
$30$ $8.78$ $8.14$
$40$ $8.26$ $6.88$
Each indicator depends on two features $f_{1}$ and $f_{2}$ and real number $l\in[0,1]$. Results are reported for neural network trained to predict eigenspaces for elliptic eigenproblem with $k_{1}=k_{2}$ and $k_{1}\neq k_{2}$ (see Appendix F.1 for description) with loss function $L_{2}(A,B;z)$. For each indicator we provide values for several $l$ averaged over $1000$ randomly selected feature pairs $f_{1}$, $f_{2}$. Results are reported in Table 7, Table 8, Table 9.
All indicators clearly demonstrate the improve in smoothness when the size of embedding $N_{\text{subspace}}$ increases. Results for average cosine indicator (50) indicate that learned mapping effectively average information about subspaces for distinct features. It becomes especially clear if one compares results for $l=1.0$ with average cosines computed between targets $\frac{1}{D}\sum_{i=1}^{D}\cos_{i}(t_{1},t_{2})$: for $k_{1}=k_{2}$ average cosine is $0.51$; for $k_{1}\neq k_{2}$ average cosine is $0.53$.
Table 8: Taylor smoothness indicator (49).
$k_{1}=k_{2}$ $k_{1}\neq k_{2}$
$N_{\text{subspace}}$ $l=0.25$ $l=0.5$ $l=0.75$ $l=1.0$ $l=0.25$ $l=0.5$ $l=0.75$ $l=1.0$
$10$ $2.41$ $5.1$ $7.48$ $10.71$ $1.8$ $3.87$ $5.99$ $8.21$
2020 $2.12$ $4.57$ $7.05$ $9.68$ $1.72$ $3.77$ $5.86$ $8.08$
$30$ $1.88$ $4.09$ $6.34$ $8.73$ $1.71$ $3.77$ $5.87$ $8.08$
$40$ $1.76$ $3.84$ $5.98$ $8.22$ $1.42$ $3.15$ $4.94$ $6.8$
Table 9: Average cosine indicator (50).
$k_{1}=k_{2}$ $k_{1}\neq k_{2}$
$N_{\text{subspace}}$ $l=0.25$ $l=0.5$ $l=0.75$ $l=1.0$ $l=0.25$ $l=0.5$ $l=0.75$ $l=1.0$
$10$ $0.81$ $0.74$ $0.67$ $0.51$ $0.82$ $0.74$ $0.67$ $0.53$
2020 $0.94$ $0.88$ $0.82$ $0.68$ $0.94$ $0.88$ $0.83$ $0.72$
$30$ $0.97$ $0.93$ $0.88$ $0.78$ $0.97$ $0.93$ $0.89$ $0.81$
$40$ $0.98$ $0.95$ $0.91$ $0.82$ $0.99$ $0.96$ $0.93$ $0.87$
F.6 Subspace regression combined with LOBPCG
Subspace regression can be also used to improve results for classical iterative eigensolvers. We demonstrate this for LOBPCG, which is a matrix-free iterative eigensolver that can approximate extremal eigenspaces [19]. A notable feature of LOBPCG is a possibility of hot start: when approximation to eigenspace of interest is available, it can be used at the initialisation to speed up convergence. In Table 10 we report such speed up for elliptic eigenproblems described in Appendix F.1. Metrics in Table 10are computed for test set, and maximal number of iterations for LOBPCG is set to $1000$. Convergence plots are available in FIgure 7(a) and Figure 7(b). It is evident that initialisation by subspace regression lead to both smaller number of iteration and better final error. Note, that the overall cost of method is dominated by the cost of iterations.
Table 10: Performance of LOBPCG with initialisation by subspace regression compared with random initialisation, $N_{\text{it}}$ number of iterations until convergence.
$k_{1}=k_{2}$ $k_{1}\neq k_{2}$
initialisation $N_{\text{subspace}}$ $N_{\text{it}}$ relative error $\pm$ std $N_{\text{it}}$ relative error $\pm$ std
subspace regression $10$ $410\pm 245$ $0.093\pm 1.109$ $332\pm 213$ $0.028\pm 0.179$
subspace regression 2020 $288\pm 197$ $0.159\pm 4.12$ $249\pm 174$ $0.02\pm 0.087$
subspace regression $30$ $274\pm 209$ $0.047\pm 0.554$ $227\pm 172$ $0.022\pm 0.119$
subspace regression $40$ $263\pm 200$ $0.131\pm 3.218$ $221\pm 160$ $0.02\pm 0.089$
random $685\pm 265$ $7.86\pm 30.16$ $609\pm 231$ $11.42\pm 35.49$
(a) $D=2$, $k_{1}=k_{2}$
(b) $D=2$, $k_{1}\neq k_{2}$
Figure 7: Convergence plots for LOBPCG with and without initialisation by subspace regression.